# Can LLMs Replace Human Evaluators? An Empirical Study of LLM-as-a-Judge in Software Engineering

### v1

@@ Abstract.
large language models, evaluation, human preference
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: the 34th ACM SIGSOFT International Symposium on Software Testing and Analysis; June 25–28, 2025; Trondheim, Norwayisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineeringccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Since BERT (Devlin et al., 2019) and GPT (Radford, 2018), pre-trained language models (PLMs) have been widely used in various natural language processing (NLP) tasks, such as machine translation and text summarization. With the scaling of PLM parameters, the concept of large language models (LLMs) has been proposed. Featuring up to hundreds of billions of parameters, LLMs emerge new capabilities absent on smaller models (Wei et al., 2022a), beyond solving simple linguistic tasks. These capabilities include, but not limited to, instruction following and multi-step reasoning, enabling LLMs to simulate human experts and achieve state-of-the-art performance in certain domains. Software engineering (SE) is one of the specialized domains that benefits from this trend. Many researchers and companies either emphasize their LLMs’ strong coding performance (OpenAI, 2023; Reid et al., 2024), or develop specialized code LLMs. For instance, DeepSeek-Coder-V2 (DeepSeek-AI et al., 2024) correctly generates code for 75.3% instructions in HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) with 236B parameters, second only to GPT-4o. Qwen2.5-Coder (Hui et al., 2024) achieves 88.4% Pass@1 on HumanEval with merely 7B parameters.
@@ 4.1.2. Response Generation
We deploy 12 recent code LLMs with different sizes from seven families shown in Table 1 from Hugging Face (Wolf et al., 2019). We generate responses using these LLMs with vLLM (Kwon et al., 2023) on a Ubuntu 20.04 server with two Intel Xeon Platinum 8276L CPUs, four NVIDIA A100-40GB GPUs, and 256 GB RAM. For each instruction, we randomly select three LLMs to respond, yielding three responses $A,B,C$. For pairwise comparisons, we create three response pairs $(A,B),(A,C),(B,C)$, and another three pairs $(B,A),(C,A),(C,B)$ in order to check if LLM-as-a-judge methods yield consistent judgment after reversing the order within a response pair. Thus, we obtain 150 responses and 300 response pairs per task, resulting in 450 responses and 900 pairs in total.
@@ 5.1. RQ1: Alignment with Human Scores
Translation Generation Summarization
Method $\rho$ $R$ $\tau$ $\rho$ $R$ $\tau$ $\rho$ $R$ $\tau$
@@ 5.2. RQ2: Score Characteristics
Translation Generation Summarization
Category $\rho_{\text{conv}}$ $\rho_{\text{other}}$ $\rho_{\text{inner}}$ $\rho_{\text{conv}}$ $\rho_{\text{other}}$ $\rho_{\text{inner}}$ $\rho_{\text{conv}}$ $\rho_{\text{other}}$ $\rho_{\text{inner}}$
@@
Translation Generation Summarization
LLM Sizes Compared $\rho_{\min}$ $\rho_{\max}$ $\rho_{\min}$ $\rho_{\max}$ $\rho_{\min}$ $\rho_{\max}$
@@ 5.3. RQ3: Pairwise Comparison versus Individual Scoring
Translation Generation Summarization
Method Acc. Agr. Acc. Agr. Acc. Agr.
@@ 6.2. Implications of Findings
    • For code summarization, LLM-as-a-judge methods should not be used directly or alone due to their low alignment with human judgments. However, with carefully designed prompts to mitigate common LLM biases, they can still serve as a valuable complement to conventional metrics.
@@
    There are still gaps to be bridged between LLM and human behavior during evaluation:
    • During evaluation, LLMs typically rely solely on the predefined evaluation criteria. In contrast, human evaluators can compare multiple responses, implicitly identifying common strengths and weaknesses to streamline the evaluation process. To bridge this gap, researchers could enhance LLM-as-a-judge methods by asking LLMs to summarize insights from previous evaluation sessions. These insights could be included in the prompt to provide additional evaluation context. Multiple responses are also valuable for LLMs to make comparisons.

### v2

@@ Abstract.
large language models, model evaluation, human preference
copyright: ccdoi: 10.1145/3728963journalyear: 2025journal: PACMSEjournalvolume: 2journalnumber: ISSTAarticle: ISSTA086publicationmonth: 7ccs: Software and its engineeringccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Since BERT (Devlin et al., 2019) and GPT (Radford, 2018), pre-trained language models (PLMs) have been widely used in various natural language processing (NLP) tasks, such as machine translation and text summarization. With the scaling of PLM parameters, the concept of large language models (LLMs) has been proposed. Featuring up to hundreds of billions of parameters, LLMs emerge new capabilities absent on smaller models (Wei et al., 2022a), beyond solving simple linguistic tasks. These capabilities include, but are not limited to, instruction following and multi-step reasoning, enabling LLMs to simulate human experts and achieve state-of-the-art performance in certain domains. Software engineering (SE) is one of the specialized domains that benefits from this trend. Many researchers and companies either emphasize their LLMs’ strong coding performance (OpenAI, 2023; Reid et al., 2024), or develop specialized code LLMs. For instance, DeepSeek-Coder-V2 (DeepSeek-AI et al., 2024) correctly generates code for 75.3% instructions in HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) with 236B parameters, second only to GPT-4o. Qwen2.5-Coder (Hui et al., 2024) achieves 88.4% Pass@1 on HumanEval with merely 7B parameters.
@@ 4.1.2. Response Generation
We deploy 12 recent code LLMs with different sizes from seven families shown in Table 1 from Hugging Face (Wolf et al., 2019). We generate responses using these LLMs with vLLM (Kwon et al., 2023) on an Ubuntu 20.04 server with two Intel Xeon Platinum 8276L CPUs, four NVIDIA A100-40GB GPUs, and 256 GB RAM. For each instruction, we randomly select three LLMs to respond, yielding three responses $A,B,C$. For pairwise comparisons, we create three response pairs $(A,B),(A,C),(B,C)$, and another three pairs $(B,A),(C,A),(C,B)$ in order to check if studied methods yield consistent judgment after reversing the order within a response pair. Thus, we obtain 150 responses and 300 response pairs per task, resulting in 450 responses and 900 pairs in total.
@@ 5.1. RQ1: Alignment with Human Scores
Method Translation Generation Summarization
$\rho$ $R$ $\tau$ $\rho$ $R$ $\tau$ $\rho$ $R$ $\tau$
@@ 5.2. RQ2: Score Characteristics
Category Translation Generation Summarization
$\rho_{\text{conv}}$ $\rho_{\text{other}}$ $\rho_{\text{inner}}$ $\rho_{\text{conv}}$ $\rho_{\text{other}}$ $\rho_{\text{inner}}$ $\rho_{\text{conv}}$ $\rho_{\text{other}}$ $\rho_{\text{inner}}$
@@
LLM Sizes Compared Translation Generation Summarization
$\rho_{\min}$ $\rho_{\max}$ $\rho_{\min}$ $\rho_{\max}$ $\rho_{\min}$ $\rho_{\max}$
@@ 5.3. RQ3: Pairwise Comparison versus Individual Scoring
Method Translation Generation Summarization
Acc. Agr. Acc. Agr. Acc. Agr.
@@ 6.2. Implications of Findings
    • For code summarization, LLM-as-a-judge methods should not be used directly or alone due to their insufficient alignment with human judgments. However, with carefully designed prompts to mitigate common LLM biases, they can still serve as a valuable complement to conventional metrics.
@@
    There are still gaps to be bridged between LLM and human evaluators:
    • During evaluation, LLMs typically rely solely on the predefined evaluation criteria. In contrast, human evaluators can compare multiple responses, implicitly identifying common strengths and weaknesses to streamline the evaluation process. To bridge this gap, researchers could enhance LLM-as-a-judge methods by asking LLMs to summarize insights from previous evaluation sessions. These insights could be included in the prompt to provide additional evaluation context. Multiple responses in the prompt are also valuable for LLMs to make comparisons.
@@
    • The underperformance of LLM-as-a-judge in evaluating code summaries reveals a critical misalignment between benchmark task definitions and LLM interpretations. In our experiments, CodeXGLUE expects concise, docstring-style summaries while LLMs default to detailed explanations due to their verbosity bias. To improve evaluation reliability, researchers should mitigate LLMs’ implicit, bias-influenced assumptions about the task to ensure they correctly understand task objectives.